OpenStudy (anonymous):

can someone please help me Find the 4th term of the sequence, an=880(0.5)n - 1. 55 176 440 110

jimthompson5910 (jim_thompson5910):

Replace n with 4 and evaluate an = 880(0.5)^(n-1) a4 = 880(0.5)^(4-1) I'll let you finish

OpenStudy (anonymous):

110?

jimthompson5910 (jim_thompson5910):

you got it